OpenStudy (anonymous):

Simplify the expression. i^313

OpenStudy (anonymous):

divide 313 by 4 and take the integer remainder

OpenStudy (anonymous):

4 goes on to 312 evenly so the remainder is 1 in this case. therefore \(i^{313}=i^1=i\) pattern is \[i^0=1\] \[i^1=i\] \[i^2=-1\] \[i^3=-i\] \[i^4=1\] and so on

OpenStudy (anonymous):

ok so when i divided 313 by 4 i got 78.25 would i round that to 8?

OpenStudy (anonymous):

it is the remainder you are looking for

OpenStudy (anonymous):

and then divide that by 4 again?

OpenStudy (anonymous):

by which i mean the integer remainder. 4 goes in to 313 78 times with a remainder of 1 so you are only concerned with the 1

OpenStudy (anonymous):

i.e. \(313=78\times 4+1\) ignore the first part, work only with the 1

OpenStudy (anonymous):

So you would divide 78 by 4?